The following shows Randall and Arts Inc.s income statement for the last two years. The company had assets of $7,050 million in the first year and $11,278 million in the second year. Common equity was equal to $3,750 million in the first year, and the company distributed 100% of its earnings out as dividends during the first and the second years. In addition, the firm did not issue new stock during either year Randall and Arts Inc. Income Statement For the Year Ending on December 31 (Millions of dollars) Year 2 Yar 1 Net Sales 3,810 1,610 191 1,801 2,009 271 1,738 695 1,043 3,000 1,495 120 1,615 1,385 Operating costs except depreciation and amortization Depreciation and amortization Total Operating Costs Operating Income (or EBIT) Less: Interest Earnings before taxes (EBT) Less : Taxes (40%) Net Income 1,274 510 764

Calculate the profitability ratios of Randall and Arts Inc. in the following table. Convert all calculations to a percentage rounded to two decimal places Ratio Value Year 2 Year 1 | Operating margin Profit margin Return on total assets Return on common equity Basic earning power 46.17% 27.38% 10.84% 20.37% 17.81% Decision makers and analysts look deeply into profitability ratios to identify trends in a companys profitability. Profitability ratios give insights into both the survivability of a company and the benefits that shareholders receive. Identify which of the following statements are true about profitability ratios. Check all that apply. A higher operating margin than the industry average indicates either lower operating costs, higher product pricing, or both. □ An increase in the return on assets ratio implies an increase in the assets a firm owns. If a companys operating margin increases but its profit margin decreases, it could mean that the company paid more in interest or taxes. If a company issues new common shares but its net income does not increase, return on common equity will increase
Answer #1

1.Operating Margin Year 2 = 2009/3810 = 52.73%
Profit Margin Year 1 = 764/3000 = 25.47%
Return on Total Assets Year 2 = Net income/Total Assets = 1043/11,278 = 9.25%
Return on Equity Year 2 = Net income/Total Equity = 1043/3750 = 27.81%
Basic Earning Power Year 1 = EBIT/Total Assets = 1385/7050 = 19.65%

2. Option a,c are correct
Option b and d are incorrect.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well